The production department is proposing the purchase of an automatic insertion machine. It has identified 3 machines and has asked the accountant to analyze them to determine the best cash payback.

Machine A Machine B Machine C
Annual cash flow $40,000 $50,000 $75,000
Average investment $300,000 $250,000 $500,000

a.Machine C

b.Machine A

c.Machine B

d.All of these choices are equal.
Answer #1

Payback period=Average investment/Annual cash flows

A B C
Payback period (300,000/40000)=7.5 years (250000/50,000)=5 years (500,000/75000)=6.67 years(Approx)

Hence B is the best machine in terms of lowest payback period.